If you could afford exactly 6 apples and 14 bananas, or 10 apples and 6 bananas, then if (2) you spent all of your income on bananas, how many bananas could you buy?
Answer #1

Answer

increase in apple=10-6=4

decrease in banana=14-6=8

The increase in 1 apple decreases two banana means the price of banana is half of the apple.

From first pair 6 apple and 14 banana and if we decrease the 6 apples then you can purchase 12 more banana. it means you can buy in total=14+12=26

from the second pair, if you decrease 10 apples, then you can buy 20 more banana so the total =6+20=26

You can buy 26 banana.